# A Cost-Benefit Analysis of On-Premise Large Language Model Deployment: Breaking Even with Commercial LLM Services

### v1

@@ Abstract
Large language models (LLMs) are becoming increasingly widespread. Organizations that want to use AI for productivity now face an important decision. They can subscribe to commercial LLM services or deploy models on their own infrastructure. Cloud services from providers such as OpenAI, Anthropic, and Google are attractive because they provide easy access to state-of-the-art models and are easy to scale. However, concerns about data privacy, the difficulty of switching service providers, and long-term operating costs have driven interest in local deployment of open-source models. This paper presents a cost-benefit analysis framework to help organizations determine when on-premise LLM deployment becomes economically viable compared to commercial subscription services. We consider the hardware requirements, operational expenses, and performance benchmarks of the latest open-source models, including Qwen, Llama, Mistral, and etc. Then we compare the total cost of deploying these models locally with the major cloud providers subscription fee. Our findings provide an estimated breakeven point based on usage levels and performance needs. These results give organizations a practical framework for planning their LLM strategies.
Index Terms:
Large Language Models, On-Premise Deployment, Cost-Benefit Analysis, Total Cost of Ownership.
@@ I Introduction
The rapid advancement of LLMs has encouraged companies and organizations to integrate them into user-facing services, which enables more responsive and adaptive interactions that cater to diverse user needs [1, 3, 2]. As adoption grows, organizations face a critical strategic choice: whether to rely on commercial cloud-based services subscription or to invest in their own on-premise deployment infrastructure [4, 5]. Through APIs and subscription services, providers like OpenAI, Anthropic, and Google are making their state-of-the-art models easy to access. However, the expenses of these solutions can rise rapidly when used by a large group of people [6]. They also bring up issues with regulatory compliance, data protection, and the challenge of transferring service providers [7, 8]. For instance, Desai et al. [9] point out that privacy concerns are a key factor slowing the adoption of LLMs in financial organizations where compliance and trust are crucial.
At the same time, organizations now have the option to deploy competitive open-source models released by the open-source community, including LLaMA [10], Mistral [11], and Qwen [12]. Recent advances in GPU hardware (NVIDIA H100[13], AMD MI300X[14]) and inference optimization frameworks (vLLM[15], NVIDIA TensorRT-LLM[16], DeepSpeed[17]) also contributed to making local deployment more feasible. As a result, it is worth for enterprises to reconsider whether building their own AI infrastructure may be a more economically viable option.
Despite an increasing interest in balancing online subscriptions with local deployment, a lack of systematic, quantitative analysis comparing the two approaches remains. Organizations need a clear framework to evaluate the break-even point when local deployment becomes cost-effective compared to commercial services.
@@
    A systematic survey of current commercial LLM pricing models and open-source alternatives suitable for local deployment.
@@
    An online playground [^1] Available at: https://v0-ai-cost-calculator.vercel.app/ where enterprise users can apply the cost-benefit framework to the latest models, explore hardware/API trade-offs, and tailor break-even analysis to their workloads.
@@ II Background and Related Work
II-A LLM Deployment Paradigms
Today, modern LLM deployment follows three paradigms: cloud, on-premise, and hybrid. Cloud services provide direct access to state-of-the-art models without pre-setup, but introduce challenges such as recurring costs, data security, and sovereignty [19, 18]. On-premise deployments, in contrast, give organizations full control over infrastructure and ensure that sensitive data remain in-house rather than being used for external model training. For domains such as healthcare, finance, and law, local deployment is often preferred due to strict security and compliance requirements [20]. However, local setups require significant upfront investment and specialized expertise. To balance these trade-offs, hybrid approaches combine local execution for sensitive data with selective cloud offloading for scalability, latency reduction, and accuracy [21].
II-B Cost Analysis Literature
Companies and organizations deploying LLMs typically weigh performance and governance, but cost consistently stands out as a top concern. Prior research into LLM economy has largely focused on four key areas. Firstly, some studies aim to reduce costs without loss of accuracy, concentrating on API pricing and usage tactics like model cascades [6, 22]. Secondly, systems-level effort emphasizes on reducing inference cost per token, achieved through optimizations such as paging-based memory management, quantization, batching, and speculative decoding [15, 23, 24, 25]. Thirdly, deployment-model research examines the financial aspects of serverless inference, provisioned throughput, and adapter-sharing for multi-tenant workloads [26, 27, 28, 29]. Finally, recent researchers have modeled the total cost of ownership (TCO). This work underscores the trade-offs between spending on cloud-based API and amortized hardware costs of on-premises local deployment. They also discussed the impact of precision formats (e.g., FP8) and energy efficiency [30, 31, 32].
Despite this progress, few frameworks truly integrate the economics of cloud and on-premise. While extensive literature already covers individual topics (e.g., cloud API costs or local hardware efficiency), a unified model to guide business decisions is still missing. This study aim to bridge this gap. We seek to embed cost analysis within a comprehensive framework that considers pricing, performance, and governance together.
II-C Open-Source Model Ecosystem
The open-source LLM ecosystem has accelecrated, with many models now achieving performance as well as commercial offerings. Meta’s LLaMA-3 family provides models ranging from 8B to 405B parameters. Outperforming or matching with the commercial giants like Claude and Gemini’s 70B and 405B in reasoning and knowledge tests [33]. Alibaba’s Qwen series reinforces this pattern. Its Qwen-3 model set new standards in mathematics, coding, and multilingual understanding while remaining openly available [34].
The quick spread of powerful models makes them easier to adopt. Organizations can now run advanced LLMs themselves, which provides more control over expenses, tailoring new features, and data security. Consequently, organizations used to depend on commercial APIs are now seriously looking at open-source alternatives for their operation needs.
@@ III-A Performance Evaluation Framework
We’ve developed a thorough performance evaluation framework to ensure our cost-benefit analysis accurately reflects real-world deployment. Our assessment criteria are:
Accuracy Metrics. Selecting the right model begins with accuracy, but accuracy in isolation is insufficient for enterprise decision-making. Organizations require evidence of model performance on tasks directly relevant to their workflows. To select models that are likely to deliver practical value, we ground our evaluation in standardized benchmarks that represent complex analytical, quantitative, and technical challenges. Specifically, we adopt GPQA (graduate-level reasoning)[59], MATH-500 (mathematical problem solving)[60], MMLU-Pro (broad multi-task language understanding)[63], and LiveCodeBench (software engineering and debugging tasks)[61]. Each benchmark captures a distinct class of enterprise need: high-level research and analysis (GPQA), computationally intensive problem solving (MATH-500), cross-domain reasoning (MMLU-Pro), and applied software development support (LiveCodeBench). We then compare accuracy across commercial systems (e.g., OpenAI, Anthropic) and leading open-source models (e.g., LLaMA, QWen-3), quantifying the performance gap and assessing whether open alternatives can feasibly replace commercial APIs. The benchmark results are reported in Table I(the data are adapted from Artificial Analysis [62]), while the performance gap between open-source and commercial models is summarized in Table VI. By tying evaluation to representative tasks and aligning results with enterprise priorities, this framework ensures that model selection is not only about raw accuracy but also about measurable contributions to organizational value, balanced against cost and deployment flexibility.
@@ III-B Model Selection Criteria
Bringing accuracy and operations together, we define the following criteria for models included in our survey:
@@
    Performance Parity: Benchmark scores within 20% of leading commercial models. This threshold reflects enterprise practice, where modest performance gaps are often outweighed by considerations of cost, security, and integration.
@@
    Deployment Feasibility: Hardware requirements suitable for typical enterprise environments
@@
    License Compatibility: Open-weight models distributed under permissive licenses that allow commercial deployment.
@@
    Community Support: Evidence of an active optimization and development ecosystem, ensuring ongoing improvements and stability.
TABLE I: Performance Benchmarks for Trending Models(data from Artificial Analysis [62])
Model Total Parameters GPQA MATH-500 LiveCodeBench MMLU-Pro
@@
GPT-5 85.4% 99.4% 66.8% 87.1%
Claude-4-Sonnet 68.3% 93.4% 44.9% 83.7%
Claude-4-Opus 70.1% 94.1% 54.2% 86.0%
Grok-4 87.7% 99.0% 81.9% 86.6%
Gemini-2.5-Pro 84.4% 96.7% 80.1% 86.2%
@@ IV Commercial LLM Pricing Models
The API subscription model charges based on the number of tokens processed (input and output), suitable for integrating LLM capabilities into custom applications. Costs can vary dramatically depending on usage patterns, batch processing, and model selection. We provide detailed cost breakdowns and usage scenarios in TableII.
TABLE II: Commercial LLM API Pricing (per 1M tokens, USD)
Provider Model Input Output
OpenAI[51] GPT-5 \$1.25 \$10.00
Anthropic[52] Claude-4 Opus \$15.00 \$75.00
Anthropic Claude-4 Sonnet \$3.00 \$15.00
xAI[53] Grok-4 \$3.00 \$15.00
Google[54] Gemini 2.5 Pro \$1.25 \$10.00
@@ V-A Capital Expenditures: The Hardware
  • Data Center Grade (e.g., NVIDIA A100[57]): Highest performance, best for large-scale, multi-user deployments. Enables running the largest models and highest throughput.
@@
TABLE III: Comparison of A100-80GB and RTX 5090-32GB GPUs
@@
Table III lists the memory, power, and approximate price for NVIDIA 5090-32GB and A100-80GB GPUs. Table IV lists the hardware resources needed to set up the service.
TABLE IV: Hardware Requirements for Deployable Open-Source Models (FP8/W8A16 on A100[57]/RTX 5090[56])
Model MoE Total Params Active Params VRAM (FP8) Hardware Deployment Tokens/sec
@@
V-B Operational Expenditures: The Hidden Costs
Local deployment does not only involve upfront hardware purchases; the running costs are also not negligible. In practice, operational expenditures come from many sources, including network, storage, backup, security, and facility overhead. In this paper, we restrict attention to the electricity cost from model execution, to keep our cost model tractable and transparent. Power usage is the dominant recurring expense and the easiest to quantify.
@@ VI Cost Model and Break-Even Analysis
We developed a quantitative model to assess the cost and feasibility of matching the throughput of leading commercial LLMs. Our model uses available benchmark, hardware, and pricing data.
@@ VI-A Cost-Performance Tradeoffs of Open-Source LLM Deployment
Table I and table VI highlights that open-weight models can deliver competitive performance. While the largest open models (e.g., Kimi-K2, GLM-4.5, Qwen3-235B) require multi-node GPU clusters costing upwards of \$200k, their accuracy on enterprise-relevant benchmarks place them within striking distance of the strongest closed models. This demonstrates that, from a purely technical standpoint, open alternatives can approach commercial state-of-the-art, albeit with greater operational complexity.
More importantly, the gap between “large” and “medium” deployments is narrower than might be expected. Medium-scale models such as gpt-oss-120B, GLM-4.5-Air, and Llama-3.3-70B can run comfortably on just two A100-80GB GPUs (\$30k), with accuracy reduction typically within 10%. The cost of ownership is thus significantly lower while still delivering practical accuracy on reasoning, coding, and domain-specific tasks.
Finally, small models (e.g., EXAONE 4.0 32B, Qwen3-30B, Magistral Small) demonstrate that even sub-30B parameter deployments are feasible on a single consumer-grade RTX 5090 (\$2k). With comparable performance to medium open models, the small models are well-suited for small and medium enterprises that prioritize cost efficiency and local control. While they fall short of the very highest benchmarks, the observed performance differences between 30B- and 70B-class models are relatively modest in practice, suggesting that smaller deployments can already cover a broad range of enterprise use cases.
While commercial APIs retain a slim lead in peak accuracy and efficiency, open-weight models now offer viable, cost-effective alternatives. The performance gap between large, medium, and small open deployments is far narrower than the order-of-magnitude differences in hardware cost. For many organizations, deploying a medium or even small open model locally provides a sustainable “break-even” alternative that balances capability, cost, and independence from external providers.
TABLE V: Hardware and Capacity Analysis for Open-Source Models (Electricity at \$0.15/kWh)
Model Hardware Cost Electricity Cost/Month Throughput (tok/sec) Token Capacity/Month
Large Open Models
Kimi-K2 \$240k \$126.72 800 506.9M
GLM-4.5 \$90k \$47.52 400 253.4M
Qwen3-235B \$60k \$31.68 400 253.4M
Medium Open Models
gpt-oss-120B \$30k \$15.84 220 139.4M
GLM-4.5-Air \$30k \$15.84 200 126.7M
Llama-3.3-70B \$15k \$7.92 190 120.4M
Small Open Models
EXAONE 4.0 32B \$2k \$13.20 200 126.7M
Qwen3-30B \$2k \$13.20 180 114.0M
Magistral Small \$2k \$13.20 150 95.0M
@@ VI-B Cost Model
The initial one-time infrastructure setup cost is defined as:
@@
Based on operational requirements for business hours (8 hours/day, 20 days/month), the monthly electricity cost for LLM inference is defined as:
@@
Given the cost of infrastructure setup and monthly electricity cost for inference, we define the local deployment cost as:
@@
where:
  • $t$: Number of months after deployment
  • $C_{\text{GPU}}$: Cost per GPU (e.g., A100-80G or RTX 5090-32G)
  • $N_{\text{GPU}}$: Number of GPUs
  • $P_{\text{GPU}}$: GPU power consumption (Watts)
  • $R_{\text{electricity}}$: Business electricity rate (\$/kWh)
  • $H_{\text{operation}}$: Operating hours per month
  • $T_{\text{throughput}}$: Token generation rate (tokens/second)
@@ VI-C Commercial API Throughput Analysis
To compare against local deployments, we normalize costs by the token generation capacity $Q_{capacity}$ achievable on open-source hardware (from Table IV).
For a commercial API, the equivalent cost of generating the same number of tokens per month is defined as:
@@
Given the cost of generating the equivalent number of toekns per month with the open-source models, we define the API usage cost as:
$C_{API}(t)=C_{API}(Q_{capacity})\times t$ (5)
where:
  • $Q_{capacity}$: Monthly token generation capacity of the open model and defined as:
  • $C_{input}$: API price for 1M input tokens
  • $C_{output}$: API price for 1M output tokens
  • $2:1$ ratio reflects typical workloads where completions are longer than prompts
Thus, for any open model with capacity $Q_{capacity}$, the equivalent API expenditure scales linearly with both token throughput and the API pricing structure. This allows direct comparison of one-time hardware investment plus electricity costs versus recurring API charges.
VI-D Break-Even Intersection Analysis
We express local and API costs as time-dependent functions to identify the operating horizon where local deployment becomes more economical.
  • $C_{local}(t)$ is a step function starting at hardware cost, growing linearly with electricity costs.
  • $C_{API}(t)$ is a linear function that grows with the usage of token usage capacity.
To achieve break-even condition, we need to find $t$ that satisfies:
$C_{local}(t)=C_{API}(t)$ (7)
Solving this equation yields the critical time $t^{*}$, the point at which the cumulative cost of local deployment equals the cumulative API expenditure. For $t>t^{*}$, on-premise deployment becomes more cost-effective, particularly under sustained high-throughput workloads. In the following section, we present the calculated break-even points $t^{*}$ across different open-source and commercial LLMs to quantify the threshold at which local hosting achieves economic advantage.
Figure 1: Break-even analysis for open-source LLMs: Large (top), Medium (middle), and Small (bottom) models.

### v2

@@ Abstract
Large language models (LLMs) are becoming increasingly widespread. Organizations that want to use AI for productivity now face an important decision. They can subscribe to commercial LLM services or deploy models on their own infrastructure. Cloud services from providers such as OpenAI, Anthropic, and Google are attractive because they provide easy access to state-of-the-art models and are easy to scale. However, concerns about data privacy, the difficulty of switching service providers, and long-term operating costs have driven interest in local deployment of open-source models. This paper presents a cost-benefit analysis framework to help organizations determine when on-premise LLM deployment becomes economically viable compared to commercial subscription services. We consider the hardware requirements, operational expenses, and performance benchmarks of the latest open-source models, including Qwen, Llama, Magistral, and etc. Then we compare the total cost of deploying these models locally with the major cloud providers subscription fee. Our findings provide an estimated breakeven point based on usage levels and performance needs. These results give organizations a practical framework for planning their LLM strategies.
@@ I Introduction
The rapid development of Large Language Models(LLMs) has driven organizations to apply them in user-facing services for more adaptive user-facing services [1, 3, 2]. As adoption grows, organizations face a critical strategic choice: whether to rely on commercial cloud-based services subscription or to invest in their own on-premise deployment infrastructure [4, 5]. Providers such as OpenAI, Anthropic, and Google offer easy API access to state-of-the-art and agentic models [6], but costs scale quickly with usage [7]. These solutions also raise concerns over compliance, data protection, and the challenge of transferring service providers [8, 9]. For example, Desai et al. [10] note that privacy issues hinder LLM adoption in finance, where trust and regulation are critical.
At the same time, organizations now have the option to deploy open-source models, including LLaMA [11], Mistral [12], and Qwen [13]. Recent advances in GPU hardware (NVIDIA H100[14], AMD MI300X[15]) and inference optimization frameworks (vLLM[16], NVIDIA TensorRT-LLM[17], DeepSpeed[18]) also contributed to making local deployment more feasible. As a result, it is worth for enterprises to reconsider whether building their own AI infrastructure may be a more economically viable option.
Despite rising interest in hybrid deployment, few systematic and quantitative comparisons of the two approaches exist. Organizations need a framework to determine when local deployment outweighs commercial services in cost-effectiveness.
@@
    A survey of commercial LLM pricing models and open-source alternatives suitable for local deployment.
@@
    A playground [^1] Available at: https://v0-ai-cost-calculator.vercel.app/ where enterprise users can apply the cost-benefit framework to explore hardware/API trade-offs.
@@ II Background and Related Work
II-A LLM Deployment and Cost Trade-offs
Modern LLM deployment spans three paradigms: cloud, on-premise, and hybrid. Cloud services provide immediate access to state-of-the-art models but raise issues of recurring cost, data security, and sovereignty [20, 19]. On-premise deployments offer full control and compliance for sensitive domains such as healthcare, finance, and law [21], yet require significant upfront investment. Hybrid solutions balance these trade-offs by running critical workloads locally while offloading scalable or latency-sensitive tasks to the cloud [22].
Prior cost analyses of LLM deployment focus on four directions: (1) reducing API expenditure through pricing strategies and model cascades [7, 23], (2) minimizing inference cost per token via quantization, batching, and speculative decoding [16, 24, 25, 26], (3) improving system efficiency through serverless or multi-tenant provisioning [27, 28, 29, 30], and (4) modeling TCO considering hardware amortization, precision formats, and energy efficiency [31, 32, 33].
Despite these advances, few frameworks jointly analyze the economics of cloud and local deployment. Existing studies typically isolate API or infrastructure costs and leave a gap in unified evaluation. This work addresses this gap by integrating pricing, performance, and governance into a single cost-benefit framework for LLM deployment.
II-B Open-Source Model Ecosystem
The open-source LLM ecosystem has accelerated, with many models achieving performance as well as commercial models. Meta’s LLaMA-3 family (8B–405B parameters) matches or surpasses models like Claude and Gemini in reasoning and knowledge tests [34]. Similarly, Alibaba’s open-source Qwen-3 sets new benchmarks in mathematics, coding, and multilingual understanding [35]. The rapid spread of such models allows organizations to run advanced LLMs in-house for greater cost control, feature customization, and data security. As a result, many formerly reliant on commercial APIs are now considering open-source alternatives.
@@ III-A Performance Evaluation Framework
We developed a performance evaluation framework to ensure our cost-benefit analysis reflects deployment conditions.
Accuracy Metrics. While accuracy is essential, enterprises require evidence of task-specific performance. We evaluate models across diverse benchmarks encompassing reasoning, mathematics, coding, and multi-domain understanding [66, 61, 62, 65, 63]. This captures analytical, computational, and applied challenges representative of enterprise workloads.
We then compare open-weight models (e.g., LLaMA, Qwen-3) with commercial models (e.g., ChatGPT, Claude) to quantify performance gaps and assess the feasibility of replacing APIs with open-source alternatives. Results are presented in Table I (adapted from [64]) and summarized in Table IV. This evaluation in realistic tasks ensures organization’s model selection aligns with their needs and budget.
@@ III-B Model Selection Criteria
Integrating accuracy and operational factors, we define four criteria for model inclusion:
@@
    Performance Parity: Benchmark scores within 20% of top commercial models, reflecting enterprise norms where small accuracy gaps are offset by cost, security, and integration benefits.
@@
    Deployment Feasibility: Hardware requirements suitable for typical enterprise infrastructure.
@@
    License Compatibility: Open-weight models under permissive licenses enabling commercial use.
@@
    Community Support: Active development communities ensuring continuous improvement and stability.
TABLE I: Performance Benchmarks and API Pricing for Leading LLMs (data from Artificial Analysis [64])
Model Total Params GPQA MATH-500 LiveCodeBench MMLU-Pro API Cost (In/Out, USD / 1M tokens)
@@
GPT-5 (by OpenAI [52]) 85.4% 99.4% 66.8% 87.1% \$1.25 / \$10.00
Claude-4 Opus (by Anthropic [53]) 70.1% 94.1% 54.2% 86.0% \$15.00 / \$75.00
Claude-4 Sonnet (by Anthropic [53]) 68.3% 93.4% 44.9% 83.7% \$3.00 / \$15.00
Grok-4 (by xAI [54]) 87.7% 99.0% 81.9% 86.6% \$3.00 / \$15.00
Gemini 2.5 Pro (by Google [55]) 84.4% 96.7% 80.1% 86.2% \$1.25 / \$10.00
@@ IV Commercial LLM Pricing Models
The API subscription model charges per processed token (input and output), making it suitable for integrating LLMs into custom applications. Costs vary by usage, batching, and model choice. Detailed breakdowns appear in Table I.
@@ V-A Capital Expenditures: The Hardware
  • Data Center Grade (e.g., NVIDIA A100[58]): Great performance for large-scale, multi-user deployments. Enables running the largest models and highest throughput.
@@
TABLE II: Comparison of A100-80GB and RTX 5090-32GB GPUs
@@
Table II lists the memory, power, and approximate price for NVIDIA 5090-32GB and A100-80GB GPUs. Table III lists the hardware resources needed to set up the service.
TABLE III: Hardware and Capacity Summary for Deployable Open-Source Models (FP8/W8A16 on A100[58]/RTX 5090[57], Electricity at \$0.15/kWh)
Model MoE VRAM (FP8) Hardware Deployment Throughput (tok/sec) Hardware Cost Token Capacity/Month
@@ VI Cost Model and Break-Even Analysis
We developed a quantitative model to assess the cost of matching the throughput of leading commercial LLMs. Our model uses available benchmark, hardware, and pricing data.
@@ VI-A Cost-Performance Tradeoffs of Open-Source LLM Deployment
Tables I and IV show that open-weight models can deliver competitive performance. Although large open models such as Kimi-K2, GLM-4.5, and Qwen3-235B require GPU clusters costing over \$200k, their accuracy on enterprise benchmarks places them close to leading closed models. This suggests that, technically, open alternatives can approach commercial state-of-the-art, albeit with higher operational complexity.
Notably, the gap between “large” and “medium” deployments is smaller than expected. Medium-scale models like gpt-oss-120B, GLM-4.5-Air, and Llama-3.3-70B run efficiently on two A100-80GB GPUs (\$30k), with less than 10% accuracy loss. They thus offer substantially lower ownership costs while maintaining strong performance across reasoning, coding, and domain-specific tasks.
Small models such as EXAONE 4.0 32B, Qwen3-30B, and Magistral Small show that sub-30B deployments are feasible on a single consumer-grade RTX 5090 (\$2k). With performance comparable to medium models, they suit small and mid-sized enterprises prioritizing cost efficiency and local control. Although they trail the highest benchmarks, the practical performance gap between 30B- and 70B-class models remains modest, indicating that smaller deployments can meet a wide range of enterprise needs.
While commercial APIs still hold a slight edge in peak accuracy and efficiency, open-weight models now provide viable, cost-effective alternatives. The performance gaps across large, medium, and small open deployments are far narrower than their order-of-magnitude hardware cost differences. For many organizations, deploying a medium or even small open model locally offers a sustainable break-even option balancing capability, cost, and autonomy from external providers.
@@ VI-B Cost Model
The one-time infrastructure cost is defined as:
@@
Assuming business operation of 8 hours/day and 20 days/month, the monthly electricity cost is:
@@
The total local deployment cost becomes:
@@
where $t$ is the number of months after deployment, $C_{\text{GPU}}$ and $N_{\text{GPU}}$ denote per-unit GPU cost and quantity, $P_{\text{GPU}}$ power consumption, $R_{\text{electricity}}$ the electricity rate, and $H_{\text{operation}}$ the monthly operating hours.
@@ VI-C Commercial API Throughput Analysis
For a fair comparison, API costs are normalized by the token generation capacity $Q_{\text{capacity}}$ achievable on local hardware:
The equivalent API cost for producing the same token volume per month is:
@@
and scales linearly with time:
$C_{\text{API}}(t)=C_{\text{API}}(Q_{\text{capacity}})\cdot t$ (6)
Here $C_{\text{input}}$ and $C_{\text{output}}$ are API prices per million input/output tokens, using a 2:1 ratio typical of real workloads.
VI-D Break-Even Analysis
Local and API costs are modeled as time-dependent functions:
$C_{\text{local}}(t)=C_{\text{API}}(t)$ (7)
Solving yields the break-even time $t^{*}$, where cumulative local and API costs are equal. For $t>t^{*}$, local deployment becomes more economical, particularly under sustained, high-throughput usage. The next section reports $t^{*}$ values across open-source and commercial models.